OpenStudy (anonymous):

express x^2 + 6x + 4 in the form (x+a)^2 -b Help please!

jigglypuff314 (jigglypuff314):

x^2 + 2xy + y^2 = (x+y)^2 and you should know that x^2 + 6x + 9 = (x+3)^2 so how would you get x^2 + 6x + 4 like that? you can add by a value that equals zero so x^2 + 6x + 4 + 5 - 5 then you would simplify it to what you need it to be x^2 + 6x + 9 - 5 which then can be factored to (x+3)^2 - 5 feel free to ask if you don't understand how that worked :)

OpenStudy (anonymous):

thank you, I'm assuming that x^2 + 2xy + y^2 = (x+y)^2 is the specific formula for this type of question?

jigglypuff314 (jigglypuff314):

it is what you should think of while solving this type of question :)

OpenStudy (anonymous):

okay thank you! :)